OpenStudy (anonymous):

will give medal for help Kelsey painted one fourth of her bedroom in three seventh of an hour. At this rate, how long would it take her to paint the entire room? three over twenty eight hour twenty eight over three hours twelve over seven hours seven over twelve hour

OpenStudy (tkhunny):

This is a proportion. \(\dfrac{1/4\;room}{3/7\;hour} = \dfrac{1\;room}{How\;Many\;Hours}\)

OpenStudy (boldjon):

Kelsey painted one-fourth of her bedroom in three-sevenths of an hour. At this rate, how long would it take her to paint the entire room? three over twenty-eight of an hour twenty-eight over three of an hour twelve over seven of an hour seven over twelve of an hour

OpenStudy (anonymous):

r*t = d this is the general form of this type of problem. however, in this case, the distance is the job (or jobs) to be done and the rate is how much of the job is cpmpleted in a certain amount of time... r = (d/t)

OpenStudy (boldjon):

1.71 which as a fraction is 12/7

OpenStudy (anonymous):

or you can set it up as a proportion... as @tkhunny did